OpenStudy (anonymous):

Graphs of functions, finding (GoF)(-3)=___

OpenStudy (amistre64):

..........

OpenStudy (amistre64):

do people not have the sense about them to see what they are posting lol

OpenStudy (anonymous):

It has a graph attach a graph, but it obviously didnt go through.

OpenStudy (amistre64):

:)

OpenStudy (amistre64):

if its a graph; go over to x = -3 and then go up or down to find the curve; then see what y value you get from it

OpenStudy (anonymous):

So, at x=-3 f(x)= 4, but the g(x) is and absolute value graph only going up to 3. This hmwk is only and I put in DNE (does not exist) but it told me it was wrong and that the answer was -1.

OpenStudy (anonymous):

This hmwk is online*

OpenStudy (amistre64):

whats the link?

OpenStudy (anonymous):

link? it's a private class online on WAMAP, I'll try to find a way for you to see it.

OpenStudy (anonymous):

did you see the attachment?

OpenStudy (amistre64):

i did :) f(x) = -x+1 g(x) = -|x| + 3 g(f(x)) = -|-x+1| +3 ; at x=-3 we get: -|--3+1|+3 -|4| + 3 -4 + 3 = -1

OpenStudy (amistre64):

that was much easier than mind reading lol

OpenStudy (anonymous):

I'm sorry, I've never had online help before. Not used to typing math stuff in the first place.

OpenStudy (anonymous):

thank you:)

OpenStudy (amistre64):

youre welcome :)